# Multi-tracer beyond linear theory

### v1

@@ 1 Introduction
In light of the large datasets provided by galaxy surveys [1, 2, 3, 4, 5, 6, 7, 8], how can we maximize the amount of information extracted from them? Attempts to extract additional information include using alternative statistics [9, 10, 11, 12, 13, 14, 15, 16] or considering field-level analysis [17]. In this work, we focus on the multi-tracer (MT) approach as a tool to extract more information from galaxy clustering. The multi-tracer method has been extensively discussed as a way to improve the measurement of primordial non-Gaussianities and redshift-space distortions [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. Moreover, MT serves as a central tool for the SPHEREx collaboration in improving their results on primordial non-Gaussianities [7, 44].
@@
Moreover, we consider for the first time a MT non-linear galaxy clustering analysis with more than two tracers. We show that, for realistic survey specifications, two tracers is the optimal number of tracers. Furthermore, we study the dependence of the MT information gain on the total tracer number density. We find that, while the MT gains are limited to the very high number density case, when modeling with non-linear bias coefficients MT overtakes ST already for number densities of $10^{-4}h^{3}$pc-3. We also consider a non-balanced split between the two subsamples, in which one of the samples is denser compared to the other. We find that most of the MT gains are present even in the case in which one of the samples correspond to $\sim 10\%$ of the total sample. This result makes easier the task of finding two samples with different bias parameters, since one can restrict to the $10\%$ less homogeneous subsample. We also discuss how MT can be used to select subsamples with different Fingers-of-God (FoG) suppression scale, such that one could use different scale cuts for different tracers. Finally, we forecast the sensitivity of current and future galaxy surveys.
@@ 2.1 The large-scale bias expansion
and their respective bias parameters $b_{\mathcal{O}}^{T}(z)$, with $b_{1}=b_{\delta}$. These operators are constructed as functions of the gravitational and velocity potentials $\Phi_{g}$ and $\Phi_{v}$, respectively. In this work, we consider only the operators in Eq. (2.2) that contribute to the one-loop power spectrum. [^1] Notice that other third-order operators, such as $\delta^{3},\delta\mathcal{G}_{2}[\Phi_{g}]$ and $\mathcal{G}_{3}[\Phi_{g}]$, in principle, contribute to the one-loop power spectrum, but they can be removed by renormalization [64]. Another central element in the bias expansion is the presence of a stochastic component $\epsilon^{T}$ [65, 66], which leads to corrections of small-scale modes onto large scales, such as the shot-noise term.
@@
where $\bar{n}_{T}$ is the number density of the tracer $T$.
@@
where $\delta^{K}_{ij}$ is the Kronecker delta. It is often assumed that $P_{\rm stoch}^{t_{i}t_{j}}(k,z)=0$ for $j\neq i$, which implies that the stochastic fields $\epsilon^{t_{i}}$ and $\epsilon^{t_{j}}$ do not correlate, as small scale processes for the tracer $i$ are, in principle, independent of those for the tracer $j$. Noise suppression can substantially enhance the information extracted out of galaxy clustering [67]. Despite the cross-noise being relatively small compared to the diagonal part [68], as predicted by the halo model [69], it has been shown that exclusion effects, satellite galaxies [70] and also nonlinearities can enhance correlations on large scales [71]. Previous MT works [45, 46] have found that the inclusion of those terms do not sensibly affect the constraints on cosmological parameters when considering non-linear theory. As we see later, this is a big difference compared to the linear MT results. We discuss the impact of including the cross-stochastic term later in this work.
@@
$P_{\rm c.t.}^{{t_{i}}{t_{j}}}(k,\mu,z)$ $=\frac{k^{2}}{k_{\rm norm}^{2}}P_{\rm L}(k,z)\left[Z^{t_{i}}_{1}(k,z)\left(c^{t_{j}}_{{\rm ct},20}(z)+c^{t_{j}}_{{\rm ct},22}(z)\mu^{2}+c^{t_{j}}_{{\rm ct},24}(z)\mu^{4}\right.\right.$ (2.11)
@@ 2.2 Fisher analysis
We also assume that all the tracers overlap within a single redshift bin. Once more, we consider a Gaussian covariance for MT, since we restrict our analysis to relatively large scales $k\leq 0.15\,h/$Mpc where the Gaussian covariance shows good agreement with the covariance measured from simulations [79]. We leave the investigation of the impact of the non-Gaussian covariance on MT for a future project. We discuss the stability of the Fisher derivatives in App. A.
@@
The Fisher is then given by Eq. (2.18) summing over the monopole, quadrupole and hexadecapole contributions. Notice that the triangle inequality of the 3j Wigner symbols bounds $\ell_{3}$ to $|l_{1}-l_{2}|\leq l_{3}\leq|l_{1}+l_{2}|$. In our analysis we consider only the first three even multipoles of the tracer power spectrum, so that the sums above are all effectively bounded.
@@ 3 The bias split and MT information
In this section, we examine how MT behaves under a direct split in the bias parameters. In Sec. 3.1, we discuss the dependence of the MT results on the difference between the final bias values of the subsamples. In Sec. 3.2, we explore how likely it is to find subsamples with different nonlinear bias coefficients. Finally, in Sec. 3.3, we analyze the MT information content and the importance of the cross-stochastic term.
@@ 3.1 The dependence on the bias difference
We start by discussing the optimal multi-tracer split to maximize the information extracted from a sample. A key advantage of working within a Fisher framework is that we can directly assess the impact on the Fisher information by smoothly varying the $\Delta b$ parameters for every bias parameter separately. Fig. 1 displays the relative errors in $\omega_{\rm cdm}$, $h$ and $A_{s}$ for both ST (black lines) and MT as a function of the bias difference $\Delta b$ [see Eq. (2.27)]. The solid lines represent the results that include non-linear theory. The $b_{1}$, $b_{\delta^{2}}$, $b_{\mathcal{G}_{2}}$ and $b_{\Gamma_{3}}$ lines correspond to a tracer split in different biases while keeping all the other parameters fixed at the ST values from Eq. (2.26). Notably, in the limit $\Delta b\to 0$ the results always match the ST case (black lines), indicating that a tracer split that does not introduce a bias difference results in no relative information gain. Fig. 1 shows that a tracer split in $\Delta b_{1},\Delta b_{\delta^{2}}$ and $\Delta b_{\mathcal{G}_{2}}$ significantly improves the relative error bars for the cosmological parameters compared to the ST analysis, with $\Delta b_{\mathcal{G}_{2}}$ yielding the greatest information gain. However, the gains are minor for a split in the third-order operator $b_{\Gamma_{3}}$. [^4] The comparison between the relative gains for each of the bias parameters as a function of $\Delta b$ is complicated, since one can always change base of the bias expansion by taking linear combinations of the operators $\mathcal{O}$. Then, also the biases change correspondingly. Instead, a more meaningful question is how hard it is to obtain a $\Delta b$ for a given operator in a given basis, which we address in Sec. 3.2. Additionally, we present extra plots in App. C, where we consider $k_{\rm max}=0.05h/$Mpc. In this large-scale regime, we observe that the error bars increase significantly. Also, a split in $b_{1}$ outperforms a split in $b_{\delta^{2}}$ and the linear term becomes dominant. This suggests that, as smaller scales are included in the analysis, the importance of splitting in non-linear operators grows. We also notice some dependence of the results on the fiducial point chosen for the bias parameters, as discussed in App. B. The qualitative results, however, remain unchanged.
@@
For comparison, the dashed lines in Fig. 1 represent the results obtained using linear theory. Naturally, the error bars are smaller in that case, as fewer parameters need to be marginalized over. However, the absence of EFT corrections may lead to biased cosmological parameters, given that we are considering relatively nonlinear scales. The blue dashed line indicates a split in $b_{1}$ within linear theory, where we see a very mild dependence on $\Delta b_{1}$. As we discuss in Sec. 3.3, MT provides improved results under a linear split only when the cross-stochastic term is neglected or for very high values for $\bar{n}^{T}$. We find that a split in $\Delta b_{1}$ yields significantly better results when non-linear corrections are included, even when accounting for cross-stochasticity. This improvement is likely to happen due to extra degeneracy breaks in the non-linear operators (see Sec. 3.3), e.g. in terms proportional to $b_{1}b_{O}$.
@@
We also show in Fig. 1 the result for a tracer split in all bias parameters simultaneously, denoted as $b_{\rm all}^{++-+}$. The four $\pm$ signs in the superscript represent the signs of $b_{1}$, $b_{\delta^{2}}$, $b_{\mathcal{G}_{2}}$, $b_{\Gamma_{3}}$ for the $t_{2}$ sample, as indicated by Eq. (2.27). Interestingly, this split does not necessarily lead to better results compared, for example, to a split in $b_{\mathcal{G}_{2}}$, particularly for $\omega_{\rm cdm}$ and $h$. This suggests that the Fisher information may be more sensitive to non-trivial parameter combinations rather than the simple variation of all bias coefficients in a single direction. In Fig. 12 of App. C, we also present results for $b_{\rm all}^{+\pm\pm\pm}$, with different sign combinations. In general, we find that splitting all four bias parameters together consistently performs better than a simple split in the linear bias $b_{1}$. From here on, we show results as a function of $b_{\rm all}^{++-+}$ (hereafter referred to simply as $b_{\rm all}$), where the sign of $b_{\mathcal{G}_{2}}$ is flipped, as a proxy of a generic tracer split. We display in Fig. 2 the monopole, quadrupole and hexadecapole spectra for the MT scenario relative to the ST case. The left panel shows the effects of a split in $b_{1}$ and the right panel illustrates the case of a split in $b_{\rm all}$, where all bias parameters are varied simultaneously by the same $\Delta b$. Notice that the monopoles are not exactly the same for $\Delta b=0$ due to the $\frac{1}{2}$ factor of difference in the number density, which enters in the stochastic part.
@@
Figure 2: Monopole, quadrupole and hexadecapole normalized by the ST spectra for a bias split in $b_{1}$ (left) and $b_{\rm all}$ (right). Different colors correspond to various $\Delta b$ value. Dotted, solid and dashed lines correspond to the auto AA, cross AB and auto BB spectra, respectively.
@@ 3.2 The bias split using the halo mass
The above results indicate a very pessimistic scenario for finding tracers with different non-linear bias parameters. However, we should keep in mind that these findings are restricted to halos that follow either the separate Universe fits of Eq. (3.1) or the LIMD relations. In a more realistic scenario, several effects can enhance the difference between samples. First, assembly bias can be very relevant, particularly for $b_{\mathcal{G}_{2}}$; for instance, [63] has shown that $b_{\mathcal{G}_{2}}$ strongly deviates from the LIMD relation, indicating the possibility of finding very strong dependence of $b_{\mathcal{G}_{2}}$ on other properties beyond the halo mass. Finding a difference $\Delta b_{\mathcal{G}_{2}}\sim 1.0$ appears to be relatively realistic considering assembly bias (see Fig. 5 of [63]). Second, realistic galaxies samples may also strongly deviate from the relations that hold for halos [85]. A more comprehensive study on the feasibility of finding two samples with distinct nonlinear biases is beyond the scope of this work, and we leave it for a future project.
@@ 3.3 The multi-tracer information content
[…]
@@
Figure 4: Relative error in $A_{s}$ as a function of the maximum mode $k_{\rm max}$ included in the analysis. The solid lines represent the full MT analysis, which includes both the cross spectra and the cross stochastic terms. When the cross-stochastic parameters are removed (dotted lines), the error bars decrease, while completely removing the cross spectra (dash-dotted lines) leads to an increase in the error bars.
@@
We start by discussing the relevance of the cross spectra. We show in Fig. 4 the relative error in $A_{s}$ as a function of the maximum $k_{\rm max}$ mode considered in the analysis, fixing the tracer number density $\bar{n}^{T}=10^{-3}\,h^{3}{\rm Mpc}^{-3}$. Similar results apply for $\omega_{\rm cdm}$ and $h$. We compare the MT and ST case in two different scenarios: one with a linear modeling and a split in $b_{1}$ and another with a nonlinear modeling and with a split in $b_{\rm all}$. The solid lines describe the full analysis, including both the cross-spectra and the cross-stochastic terms. In contrast, the dashed-dotted lines correspond to a case where we removed the cross-spectra between the two tracers, keeping only the auto correlations in Eq. (2.18). For the EFT case, we observe that removing the cross spectra leads to a loss of information, with the result approaching the ST case. However, the gains are not completely lost when removing the cross-spectra, indicating that MT still overcomes ST even when considering the auto spectra alone. For the linear case, we notice that the MT full-analysis line closely follows the ST case. Moreover, removing the cross-spectra results in no information loss, as the two solid and dashed-dotted lines completely overlap.
@@
Another key question in MT analysis is whether cross-stochasticity in the cross-spectra can be neglected. [^6] Dropping of cross-stochasticity is often justified by assuming that small-scale processes of the two tracers do not correlate. However, this assumption was shown to fail due to non-linear clustering and exclusion effects [71]. An alternative to completely removing this term is to introduce simulation-based priors on this cross-stochasticity [86]. As discussed earlier, the linear case shown in Fig. 4 suggests that MT provides almost no improvement over ST when cross-stochasticity is included in the cross spectra [Eq. (2.9) for $i\neq j$]. For comparison, we display in dotted lines the same scenario but with the cross-stochastic contribution removed. In this case, a considerable information gain is observed. This indicates that the success of MT in linear theory seems to strongly depend on the assumption of no cross-stochasticity between the tracers. In the nonlinear case, we also observe an increase in information gain when removing the cross-stochastic parameters. However, different than the linear case, MT already shows a significant improvement over ST even when cross-stochasticity is included. This highlights a fundamental difference in the information budget of MT between linear and nonlinear theory, aligning with the findings of [45, 46].
@@
Figure 5: Correlation matrices between the parameters for the ST case (top), and MT both with (left) and without (right) the cross-stochastic operator. For the MT, we fix $\Delta b_{\rm all}=0.6$. To avoid cluttering the MT matrices with labels, we denote the bias, counter-terms, and stochastic parameters by numbers, arranged in the same order as in the ST case (i.e., AA, AB, then BB). On the top right, the parameter correlation score defined in Eq. (3.9) as a function of $\Delta b$.
@@
between the parameters $i$ and $j$. We present results for the ST case (top left) and for MT scenarios with a split in $b_{\rm all}$, both including (bottom left) and excluding the cross stochastic term (bottom right). Overall, we observe that cross correlations between cosmological and other parameters tend to be smaller in the MT case, particularly when cross-stochasticity is removed.

### v2

@@ 1 Introduction
In light of the large datasets provided by galaxy surveys [1, 2, 3, 4, 5, 6, 7, 8], how can we maximize the amount of information extracted from them? Attempts to extract additional information include using alternative statistics [9, 10, 11, 12, 13, 14, 15, 16] or considering field-level analysis [17]. In this work, we focus on the multi-tracer (MT) approach as a tool to extract more information from galaxy clustering. The multi-tracer method has been extensively discussed as a way to improve the measurement of primordial non-Gaussianities and redshift-space distortions [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. Moreover, MT serves as a central tool for the SPHEREx collaboration in improving their results on primordial non-Gaussianities [7, 45].
@@
Moreover, we consider for the first time a MT non-linear galaxy clustering analysis with more than two tracers. We show that, for realistic survey specifications, two tracers is the optimal number of tracers. Furthermore, we study the dependence of the MT information gain on the total tracer number density. We find that, while the MT gains are limited to the very high number density case, when modeling with non-linear bias coefficients MT overtakes ST already for number densities of $10^{-4}h^{3}$Mpc-3. We also consider a non-balanced split between the two subsamples, in which one of the samples is denser compared to the other. We find that most of the MT gains are present even in the case in which one of the samples correspond to $\sim 10\%$ of the total sample. This result makes easier the task of finding two samples with different bias parameters, since one can restrict to the $10\%$ less homogeneous subsample. We also discuss how MT can be used to select subsamples with different Fingers-of-God (FoG) suppression scale, such that one could use different scale cuts for different tracers. Finally, we forecast the sensitivity of current and future galaxy surveys.
@@ 2.1 The large-scale bias expansion
and their respective bias parameters $b_{\mathcal{O}}^{T}(z)$, with $b_{1}=b_{\delta}$. These operators are constructed as functions of the gravitational and velocity potentials $\Phi_{g}$ and $\Phi_{v}$, respectively. In this work, we consider only the operators in Eq. (2.2) that contribute to the one-loop power spectrum. [^1] Notice that other third-order operators, such as $\delta^{3},\delta\mathcal{G}_{2}[\Phi_{g}]$ and $\mathcal{G}_{3}[\Phi_{g}]$, in principle, contribute to the one-loop power spectrum, but they can be removed by renormalization [65]. Another central element in the bias expansion is the presence of a stochastic component $\epsilon^{T}$ [66, 67, 68, 69], which leads to corrections of small-scale modes onto large scales, such as the shot-noise term.
@@
where $\bar{n}_{T}$ is the number density of the tracer $T$. For the linear case, we only keep the constant shot-noise term $c_{0}$.
@@
where $\delta^{K}_{ij}$ is the Kronecker delta. It is often assumed that $P_{\rm stoch}^{t_{i}t_{j}}(k,z)=0$ for $j\neq i$, which implies that the stochastic fields $\epsilon^{t_{i}}$ and $\epsilon^{t_{j}}$ do not correlate, as small scale processes for the tracer $i$ are, in principle, independent of those for the tracer $j$. Noise suppression can substantially enhance the information extracted out of galaxy clustering [70]. Despite the cross-noise being relatively small compared to the diagonal part [71], as predicted by the halo model [72], it has been shown that exclusion effects, satellite galaxies [73] and also nonlinearities can enhance correlations on large scales [74]. While the cross-stochastic terms prevent the information gain in linear regime [75] (see also [20]), [46, 47] have found that the inclusion of those terms do not sensibly affect the constraints on cosmological parameters when considering non-linear theory. As we see later, this is a big difference compared to the linear MT results. We discuss the impact of including the cross-stochastic term later in this work.
@@
[…]
@@ 2.2 Fisher analysis
We also assume that all the tracers overlap within a single redshift bin, taken for simplicity to be at redshift z=0. Once more, we consider a Gaussian covariance for MT, since we restrict our analysis to relatively large scales $k\leq 0.15\,h/$Mpc where the Gaussian covariance shows good agreement with the covariance measured from simulations [82]. Moreover, since we focus on the gains of MT over ST and not in the overall gain by moving towards smaller scales, we work on the assumption that non-Gaussian corrections to the covariance affect the signal-to-noise of ST and MT similarly. We leave the investigation of the extend to which this assumption is valid and the impact of the non-Gaussian covariance on MT, in the same lines as [81], for a future project. Furthermore, we discuss the stability of the Fisher derivatives in App. A.
@@
The Fisher is then given by Eq. (2.18) summing over the monopole, quadrupole and hexadecapole contributions. Notice that the triangle inequality of the 3j Wigner symbols bounds $\ell_{3}$ to $|l_{1}-l_{2}|\leq l_{3}\leq|l_{1}+l_{2}|$. In our analysis we consider only the first three even multipoles of the tracer power spectrum, so that the sums above are all effectively bounded. We do not include light cone effects and the Alcock-Paczynski distortion in the analysis, since we work at a constant redshift of $z=0$.
@@ 3 The bias split and MT information
In this section, we examine how MT behaves under a direct split in the bias parameters. In Sec. 3.1, we discuss the dependence of the MT results on the difference between the final bias values of the subsamples. In Sec. 3.2, we explore how likely it is to find subsamples with different inear bias coefficients. Finally, in Sec. 3.3, we analyze the MT information content and the importance of the cross-stochastic term.
@@ 3.1 The dependence on the bias difference
We start by discussing the optimal multi-tracer split to maximize the information extracted from a sample. A key advantage of working within a Fisher framework is that we can directly assess the impact on the Fisher information by smoothly varying the $\Delta b$ parameters for every bias parameter separately. Fig. 1 displays the relative errors in $\omega_{\rm cdm}$, $h$ and $A_{s}$ for both ST (black lines) and MT as a function of the bias difference $\Delta b$ [see Eq. (2.27)]. The solid lines represent the results that include non-linear theory. The $b_{1}$, $b_{\delta^{2}}$, $b_{\mathcal{G}_{2}}$ and $b_{\Gamma_{3}}$ lines correspond to a tracer split in different biases while keeping all the other parameters fixed at the ST values from Eq. (2.26). Notably, in the limit $\Delta b\to 0$ the results always match the ST case (black lines), indicating that a tracer split that does not introduce a bias difference results in no relative information gain. Fig. 1 shows that a tracer split in $\Delta b_{1},\Delta b_{\delta^{2}}$ and $\Delta b_{\mathcal{G}_{2}}$ significantly improves the relative error bars for the cosmological parameters compared to the ST analysis, with $\Delta b_{\mathcal{G}_{2}}$ yielding the greatest information gain. However, the gains are minor for a split in the third-order operator $b_{\Gamma_{3}}$. [^4] The comparison between the relative gains for each of the bias parameters as a function of $\Delta b$ is complicated, since one can always change base of the bias expansion by taking linear combinations of the operators $\mathcal{O}$. Then, also the biases change correspondingly. Instead, a more meaningful question is how hard it is to obtain a $\Delta b$ for a given operator in a given basis, which we address in Sec. 3.2. Additionally, we present extra plots in App. C, where we analyze the dependence on $k_{\rm max}$. In the large-scale regime, we observe that the error bars increase significantly. Also, a split in $b_{1}$ outperforms a split in $b_{\delta^{2}}$ or $b_{\mathcal{G}_{2}}$ since the linear term becomes dominant. This suggests that, as smaller scales are included in the analysis, the importance of splitting in non-linear operators grows. We also notice some dependence of the results on the fiducial point chosen for the bias parameters, as discussed in App. B. The qualitative results, however, remain unchanged.
@@
For comparison, the dashed lines in Fig. 1 represent the results obtained using linear theory. Naturally, the error bars are smaller in that case, as fewer parameters need to be marginalized over. However, the absence of EFT corrections may lead to biased cosmological parameters, given that we are considering relatively nonlinear scales. The blue dashed line indicates a split in $b_{1}$ within linear theory, where we see a very mild dependence on $\Delta b_{1}$. As we discuss in Sec. 3.3, MT provides improved results under a linear split only when the cross-stochastic term is neglected or for very high values for $\bar{n}^{T}$. We find that a split in $\Delta b_{1}$ yields significantly smaller $\sigma_{\rm MT}/\sigma_{\rm ST}$ when non-linear corrections are included (both for ST and MT), even when accounting for cross-stochasticity. This improvement is likely to happen due to extra degeneracy breaks in the non-linear operators (see Sec. 3.3), e.g. in terms proportional to $b_{1}b_{O}$.
@@
So far we have only discussed MT splits in orthogonal directions in the bias parameter space, namely changing one bias parameter at a time. However, a realistic tracer split based on a physical motivated feature is going to generate a split in all the linear and non-linear bias parameters. To account for this scenario, we introduced a homogeneous tracer split in all bias parameters simultaneously, denoted in Fig. 1 as $b_{\rm all}^{++-+}$. The four $\pm$ signs in the superscript represent the signs of $b_{1}$, $b_{\delta^{2}}$, $b_{\mathcal{G}_{2}}$, $b_{\Gamma_{3}}$ for the $t_{2}$ sample, as indicated by Eq. (2.27). Interestingly, this split does not necessarily lead to better results compared, for example, to a split in $b_{\mathcal{G}_{2}}$, particularly for $\omega_{\rm cdm}$ and $h$. This suggests that the Fisher information may be more sensitive to non-trivial parameter combinations rather than the simple variation of all bias coefficients in a single direction. In Fig. 14 of App. C, we also present results for ${\Delta}b_{\rm all}^{+\pm\pm\pm}$, with different sign combinations. In general, we find that splitting all four bias parameters together consistently performs better than a simple split in the linear bias $b_{1}$. From here on, we show results as a function of ${\Delta}b_{\rm all}^{++-+}$ (hereafter referred to simply as ${\Delta}b_{\rm all}$), where the sign of $b_{\mathcal{G}_{2}}$ is flipped, as a proxy of a generic tracer split. [^5] The optimal combination of bias parameters that maximizes the MT results is a parameter-dependent problem. Note that the sign convention for the bias (e.g., having a minus sign for $b_{\mathcal{G}_{2}}$) is also arbitrary, since linear combinations of operators can be taken to form a new basis. Also note that the definition of the operators depend on the chosen cutoff scale for the integrals, as well as on the renormalization scheme [68]. The optimal combination of bias parameters to maximize MT performance for a given parameter is to be explored in future work and the choice of the minus sign for $b_{\mathcal{G}_{2}}$ is related to the fact that it performs on average for $\omega_{\rm cdm}$ and $h$, and reasonably well for $A_{s}$ (see Fig. 14). We display in Fig. 2 the monopole, quadrupole and hexadecapole spectra for the MT scenario relative to the ST case. The left panel shows the effects of a split in $b_{1}$, for the middle panel we consider a split in $b_{\mathcal{G}_{2}}$ and the right panel illustrates the case of a split in $b_{\rm all}$, where all bias parameters are varied simultaneously by the same $\Delta b$. Notice that the MT monopole is not exactly the same as the ST monopole for $\Delta b=0$ due to the $\frac{1}{2}$ factor of difference in the number density, which enters in the stochastic part, cf. Eq. (2.12). While $b_{1}$ is responsible for a major vertical shift in the spectra, the split in $b_{\mathcal{G}_{2}}$ leads to a scale-dependent feature that becomes important when analyzing the broadband spectra. It is also understandable that the impact of $b_{\mathcal{G}_{2}}$ is larger than a split in $b_{\delta^{2}}$ and $b_{\Gamma_{3}}$. While the latter only starts to contribute at third order in perturbation theory, $b_{\delta^{2}}$ has its third-order contribution, i.e. its contraction with $\delta^{(1)}$, removed by renormalization (see Eq. 4.11 of [84]). Therefore, $b_{\mathcal{G}_{2}}$ is the only non-linear bias operator that contributes both at second and third order.
@@
Figure 2: Monopole, quadrupole and hexadecapole normalized by the ST spectra for a bias split in $b_{1}$ (left), $b_{\mathcal{G}_{2}}$ (middle) and $b_{\rm all}$ (right). Different colors correspond to various $\Delta b$ value. Dotted, solid and dashed lines correspond to the auto AA, cross AB and auto $B$ spectra, respectively.
@@ 3.2 The bias split using the halo mass
The above results indicate a very pessimistic scenario for finding tracers with different non-linear bias parameters. However, we should keep in mind that these findings are restricted to halos that follow either the separate Universe fits of Eq. (3.1) or the LIMD relations. In a more realistic scenario, several effects can enhance the difference between samples. First, assembly bias can be very relevant, particularly for $b_{\mathcal{G}_{2}}$; for instance, [64] has shown that $b_{\mathcal{G}_{2}}$ strongly deviates from the LIMD relation, indicating the possibility of finding very strong dependence of $b_{\mathcal{G}_{2}}$ on other properties beyond the halo mass. Thanks to assembly bias, finding a difference $\Delta b_{\mathcal{G}_{2}}\sim 1.0$ appears to be realistic, at least in the situation of samples with large $b_{1}$ (or correspondingly large halo masses), see Fig. 5 of [64].
We highlight that [64] considered halo masses $\sim 10^{14}\,h^{-1}\,M_{\odot}$, larger e.g. than target values for DESI [89], and therefore with large $b_{1}$. The feasibility of identifying samples with very distinct non-linear bias for low-$b_{1}$ cases still requires further investigation. Second, realistic galaxies samples may also strongly deviate from the relations that hold for halos [90]. A more comprehensive study on the feasibility of finding two samples with distinct nonlinear biases is beyond the scope of this work, and we leave it for a future project. Finally, we note that assembly bias can also be used to break the degeneracy between $b_{1}$ and the non-Gaussian bias parameter $b_{\phi}$, improving the MT constraints in $f_{\rm NL}$ [91, 26, 27].
@@ 3.3 The multi-tracer information content
$N_{\rm Tracers}\times N_{\rm bias}$ $\quad\textrm{bias parameters}\,$, (3.5)
$N_{\rm Tracers}\times N_{\rm ct}$ $\quad\textrm{counter-terms}\,$, (3.6)
$\frac{N_{\rm Tracers}\times(N_{\rm Tracers}+1)}{2}\times N_{\rm stoch}$ $\quad\textrm{stochastic parameters}\,.$ (3.7)
@@
Figure 4: Relative error in $\omega_{\rm cdm}$, $h$ and $A_{s}$ as a function of the maximum mode $k_{\rm max}$ included in the analysis. The solid lines represent the full MT analysis, which includes both the cross spectra and the cross stochastic terms. When the cross-stochastic parameters are removed (dotted lines), the error bars decrease, while completely removing the cross spectra (dash-dotted lines) leads to an increase in the error bars. We fix either $\Delta b_{1}=0.6$ or $\Delta b_{\rm all}=0.6$.
@@
We start by discussing the relevance of the cross spectra. We show in Fig. 4 the relative error in $\omega_{\rm cdm}$, $h$ and $A_{s}$ as a function of the maximum $k_{\rm max}$ mode considered in the analysis, fixing the tracer number density $\bar{n}^{T}=10^{-3}\,h^{3}{\rm Mpc}^{-3}$. We compare the MT and ST case in two different scenarios: one with a linear modeling and a split in $b_{1}$ and another with a nonlinear modeling and with a split in $b_{\rm all}$. The solid lines describe the full analysis, including both the cross-spectra and the cross-stochastic terms. In contrast, the dashed-dotted lines correspond to a case where we removed the cross-spectra between the two tracers, keeping only the auto correlations in Eq. (2.18). For the EFT case, we observe that removing the cross spectra leads to a loss of information, with the result approaching the ST case. However, the gains are not completely lost when removing the cross-spectra, indicating that MT still overcomes ST even when considering the auto spectra alone. For the linear case, we notice that the MT full-analysis line closely follows the ST case. Moreover, removing the cross-spectra results in no information loss, as the two solid and dashed-dotted lines completely overlap.
@@
Another key question in MT analysis is whether cross-stochasticity in the cross-spectra can be neglected. [^7] Dropping of cross-stochasticity is often justified by assuming that small-scale processes of the two tracers do not correlate. However, this assumption was shown to fail due to non-linear clustering and exclusion effects [74]. An alternative to completely removing this term is to introduce simulation-based priors on this cross-stochasticity [92]. As discussed earlier, the linear case shown in Fig. 4 suggests that MT provides almost no improvement over ST when cross-stochasticity is included in the cross spectra [Eq. (2.9) for $i\neq j$]. For comparison, we display in dotted lines the same scenario but with the cross-stochastic contribution removed. In this case, a considerable information gain is observed. This indicates that the success of MT in linear theory seems to strongly depend on the assumption of no cross-stochasticity between the tracers, as already pointed out by other works [75, 20]. As we discuss later in this section, the removal of the cross-stochastic leads to strong breaking of degeneracies between parameters via the cross-spectra information, e.g., by fixing the $b_{1}^{t_{1}}b_{1}^{t_{2}}$ product. When the cross-stochastic term is included, this information gain is essentially lost. In the nonlinear case, we also observe an increase in information gain when removing the cross-stochastic parameters. However, different than the linear case, MT already shows a significant improvement over ST even when cross-stochasticity is included. Extra information is therefore provided by the full $k$-dependence of the cross-spectra, allowing for a broader breaking of degeneracies among $b_{\mathcal{O}_{\alpha}}^{t_{1}}b_{\mathcal{O}_{\beta}}^{t_{2}}$, beyond the simple $b_{1}^{t_{1}}b_{1}^{t_{2}}$ that holds in the linear case. This highlights a fundamental difference in the information budget of MT between linear and nonlinear theory, aligning with the findings of [46, 47].
@@
Figure 5: Correlation matrices between the parameters for the ST case (top left), and MT both without (top right, with $\Delta b_{\rm all}=0.6$) and with (bottom left with $\Delta b_{\rm all}=0.0$ and bottom right with $\Delta b_{\rm all}=0.6$) the cross-stochastic operator. To avoid cluttering the MT matrices with labels, we denote the bias, counter-terms, and stochastic parameters by numbers, arranged in the same order as in the ST case (i.e., $b_{1}^{t_{1}},b_{1}^{t_{2}},b_{\delta^{2}}^{t_{1}},b_{\delta^{2}}^{t_{2}},\dots,c_{\rm st,22}^{t_{1}t_{1}},c_{\rm st,22}^{t_{1}t_{2}},c_{\rm st,22}^{t_{2}t_{2}}$, with the stochastic terms in the end).
@@
between the parameters $i$ and $j$. We present results for the ST case (top left) and for MT scenarios with a split in $b_{\rm all}$, both including (bottom) and excluding (top right) the cross stochastic term. Overall, we observe that cross correlations between cosmological and other parameters tend to be smaller in the MT case, particularly when cross-stochasticity is removed. For the case including the cross-stochastic contribution, we show $\Delta b_{\rm all}=0$ (equivalent to ST in information) in the bottom left and $\Delta b_{\rm all}=0.6$ on the right. The latter is clearly more diagonal, strongly indicating that MT provides a consistent way to break correlations between bias and cosmological parameters.